# Learning Successor Features with Distributed Hebbian Temporal Memory

### v2

@@ 3.2 Neural Implementation
Initially, all factor entries are zero, meaning cells have no segments. As learning proceeds, new non-zero connections grouped into segments are grown. In equation 6 we benefit from having sparse factor value vectors because its complexity depends linearly on the amount of non-zero components. And that’s usually the case in our model due to one-step Monte-Carlo learning and specific form of emission factors $F^{k}_{e}$:
$F^{k}_{e}(h^{k}_{t},o^{k}_{t})=\mathbb{I}[h^{k}_{t}\in\mathrm{col}(\varphi^{k}_{t})],$ (7)
where $\mathbb{I}$—indicator function, $\mathrm{col}(\varphi^{k}_{t})$ is a set of hidden states connected to the feature state $\varphi^{k}_{t}$ that forms a column. The form of emission factor is inspired by presumably columnar structure of the neocortex and was shown to induce sparse transition matrix in HMM (George et al., 2021).
Segment likelihood $L_{l}$, resulting from the sum-product algorithm, is calculated as if presynaptic cells are independent. However, it’s not usually the case for sparse factors. To take into account, approximately, their interdependence, we substitute the following equation for segment log-likelihood:
$\log L_{l}$ $=w_{l}\log\frac{1}{n_{l}}\sum_{u\in\mathrm{rec}(l)}w_{ul}m_{u}+$
$+\sum_{u\in\mathrm{rec}(l)}(1-w_{ul})\log m_{u},$ (8)
where $w_{pl}$—synapse efficacy or neuron specificity for segment, such that $w_{ul}=p(s_{l}=1|c^{u}_{t-1}=1)$, $w_{l}=\frac{1}{n_{l}}\sum_{u}w_{ul}$—average synapse efficacy for segment $l$, and $n_{l}$-number of cells in segment’s receptive field.
The idea that underlies the formula is to approximate between two extreme cases:
  • $p(s_{l}=1|c^{u}_{t-1}=1)\to 1$ for all $u$, which means that all cells in the receptive field are dependent and are part of one cluster, i.e., they fire together. In that case, it should be $p(s_{l})=m_{u}$ for any $u$, but we also reduce prediction variance by averaging between different $u$.
  • $p(s_{l}=1|c^{u}_{t-1}=1)\to 0$ for all $u$ means that presynaptic cells don’t form a cluster. In that case, segment activation probability is just a product of the activation probability of each cell.
@@
where $E_{l}=\log f_{l}+\log L_{l}$, $\mathrm{cells}[H^{k}_{t}]$—indexes of cells that represent states for $H^{k}_{t}$ variable. Here, we also approximate logarithmic sum with $\mathrm{max}$ operation inspired by the neurophysiological model of segment aggregation by cell (Stuart and Spruston, 2015).
The next step after computing $p(h^{k}_{t})$ distribution parameters is to incorporate information about current observations $p(h^{k}_{t}\ |\ o^{k}_{t})\propto p(h^{k}_{t})\mathbb{I}[h^{k}_{t}\in\mathrm{col}(o^{k}_{t})]$. After that, the learning step is performed. The step for closing the loop of our TM algorithm is to assign the posterior for the current step $p(h^{k}_{t}\ |\ o^{k}_{t})$ to $M^{i}_{t-1}$.
DHTM learns $f_{l}$ and $w_{ul}$ weights by Monte-Carlo Hebbian-like updates. First, $h^{i}_{t-1}$ and $h^{k}_{t}$ are sampled from their posterior distributions: $p(h^{i}_{t-1}\ |\ o^{i}_{t-1})\propto M^{i}_{t-1}$ and $p(h^{k}_{t}\ |\ o^{k}_{t})$ correspondingly. Then $f_{l}$ is updated according to the segment’s $s_{l}$ and its cell’s $c^{j}_{t}$ activity so that $f_{l}$ is proportional to several coincidences $s_{l}=c^{j}_{t}=1$ during the recent past, i.e., cell and its segment are active at the same time step:
@@
where $\alpha\in[0,1)$ is segment’s learning rate. That is, $f_{l}$ is trained in that way that it is equal to the average cell’s activation frequency, given the segment is active. It’s similar to Baum-Welch’s update rule (Baum et al., 1970) for the transition matrix in HMM, which, in effect, counts transitions from one state to another, but, in our case, the previous state (context) is represented by a group of RVs, not just one hidden RV.
Weights $w_{ul}$ are also updated by the Hebbian rule to reflect the specificity of a presynaptic $u$ for activating a segment $l$. That is, they are targeted to represent probability $p(s_{l}=1\ |\ c^{u}_{t-1}=1)$ that segment $s_{l}$ is active, given cell $u$ was active at the previous time-step. We could learn it by counting activation coincidences and mismatches. But in our algorithm it is approximated as exponential moving average of segment’s $s_{l}$ frequency activation, given $c^{u}_{t-1}=1$: $\Delta w_{ul}=\beta\cdot\mathbb{I}[c^{u}_{t-1}=1]\cdot(\mathbb{I}[s_{l}=1]-w_{ul})$, where $\beta\in[0,1)$ — learning rate.
@@ 3.3 Agent Architecture
We incorporate DHTM as a part of an RL agent. We use the same agent for other memories tested as well. The agent consists of a memory model and a feature reward function. The memory model generates SF by predicting cumulative future distributions of feature variables $\Phi^{k}$ according to equation 12. Here, we assume that feature representations of observations are good enough to define the reward function $r(\varphi^{k}_{t})$, which is also learned during interaction with the environment and, combined with SF representations, is used to estimate the action value function. For more details on employed value function decomposition, see Appendix B.
@@
The agent training procedure is outlined in Algorithm 1. For each episode, the memory state is reset to a fixed initial message with RESET_MEMORY() and action variable is set to a fixed initial action. An observation image returned by an environment (obs) is first preprocessed to get events, mimicking a simple event-based camera with a floating threshold determined from the average difference between the current and previous step image intensities. The resulting events are encoded to SDRs with a biologically inspired spatial pooling encoder described in Appendix A. In OBSERVE() routine, the memory learns to predict next feature states as described in Section 3. An agent learns associations to feature states and rewards in REINFORCE() function:
@@
We include actions in the model by forcing some of the hidden variables $H^{k}_{t}$ to represent actions. That is, we assume that information about action is included in the hidden state of the model. For example, if we have 4 actions, we set 4 states for one of the hidden variables and set its state from observation of the action.
In SAMPLE_ACION(), for each action and current belief state, TM predicts the next belief states for which SFs of horizon $T$ are formed using uniform strategy, i.e. setting action variable to uniform distribution. In most experiments, $T$ is dynamically determined from $\Phi^{k}_{t+l}$ distribution. If the distribution of $\Phi^{k}_{t+l}$ is close to uniform or states with high positive reward have high probability, the prediction cycle is stopped, rendering the prediction process similar to breadth-first search.
$\mathrm{SF}_{t+T}(\varphi^{k}=j\ |\ o_{0:t},a_{t})=\\ =\sum^{T}_{l=0}\gamma^{l}\sum_{i}p(\varphi^{k}_{t+l+1}=j\ |\ h^{k}_{t}=i)p(h^{k}_{t}=i\ |\ o_{0:t},a_{t})$ (12)
$Q(o_{0:t},a_{t})=\sum_{jk}\mathrm{SF}_{t+T}(\varphi^{k}=j\ |\ o_{0:t},a_{t})r^{k}_{j},$ (13)
where $o_{0:t}$—observation history.
@@ 4 Experiments
We test our model in a reinforcement learning task in Gridworld environment, where successor features are easy to interpret and the encoder block (see Sec. 3) is not required, and in a more challenging AnimalAI 3D environment. This section shows how different memory models affect SF learning and an RL agent’s adaptability. In our work, we compare the proposed DHTM model with LSTM (Hochreiter and Schmidhuber, 1997) and CSCG (George et al., 2021). For each setup, the results are averaged over at least five experiments with different seed values and standard deviation is shown. See Appendix D for details about baselines, their model parameters and training regimes. In Appendix E, environments and setups are more thoroughly described. The source code of experiments and algorithms’ implementations is available at https://anonymous.4open.science/r/dhtm-D02F.
@@ 4.1 Gridworld
First we test DHTM in MDP setup, that is, observation state corresponds to position in 5x5 maze. Each trial consists of two phases. In the first phase, lasting 500 episodes, there are no obstacles on the way to the terminal state that gives a reward to the agent. In the second phase, the terminal state is surrounded by a wall, so that the agent has to detour in order to get the reward, which remains in the same position. Each episode starts with the agent in the same position and ends when terminal state is entered (i.e. the reward is collected) or maximum action steps reached. Each episode, we measure the amount of action steps for agent required to reach the goal.
The results show (Fig. 3) that the agent that uses DHTM for SR formation faster converges to the optimal trajectory than CSCG. CSCG is able to converge to optimal trajectory only if the initial training dataset is gathered using random strategy after environmental changes. It requires an oracle that tells when changes occurred to gather a new training dataset. Otherwise, the agent with CSCG memory converges to suboptimal trajectories as the training batch gets very correlated trajectories following the current agent’s policy, which causes overfitting. Here, we also provide results of Q-table and SR-table agents for reference.
@@
Figure 3: Amount of steps required for the agent to reach the goal in fully observable 5x5 Gridworld using different types of memory: DHTM and CSCG. CSCG is trained in two modes. The first mode is with oracle, which tells the agent to gather some trajectories using uniform policy after environment changes. The second mode is without oracle, so the agent continues using its current strategy to gather new trajectories. At the 500th episode, the terminal state is blocked by a wall, so the agent needs to find a new way to the goal. There are also plots of learning curves for Q-table and SR-table agents for reference.
The second test was conducted in partially observable Gridworld in a 5x5 changing maze with a stationary terminal state as before. However, now all the agent can see is the colour of the cell in which it is located. The agent is also able to see the colour of an obstacle it encounters, receiving additional punishment.
As can be seen from Figure 4, in contrast to DHTM, CSCG and LSTM are not able to converge after reward is blocked at the 300th episode. Even with an oracle signalling about changes in the environment, which enables to gather new fully random trajectories, neither CSCG nor LSTM are able to relearn maze’s structure in online case scenario. It can be seen, however, that DHTM converges to a suboptimal trajectory of length 13 on average vs 9 steps for optimal trajectory, after reward is blocked. This may be the result of inferior generalisation properties of DHTM, which we discuss in Section 5.
Figure 4: Learning curves measured in action steps until the goal for the agent using different memory types for SF formation in partially observable 5x5 Gridworld environment. In the 300th episode, the goal is blocked by a wall, so the agent needs to find a new optimal trajectory. CSCG is trained in two modes. The first mode is with an oracle, which tells the agent to gather some trajectories using uniform policy after the environment changes. The second mode is without an oracle, so the agent continues using its current strategy to gather new trajectories.
In these experiments, we aimed to show that in cases of quite limited and non-stationary trajectory buffer size, an RL agent armed with DHTM memory is more adaptive than those using LSTM and CSCG. However, we do not claim that it is impossible, in principle, to train LSTM and CSCG in such a setup. We only show that DHTM notoriously fits this kind of task without numerous workarounds that would be required by those classical methods.
@@ 4.2 AnimalAI
AnimalAI is a 3D environment that is able to give an agent rich visual input in the form of RGB images. The objective of the experiment described here is to give a hint that DHTM is also able to deal with complex visual input due to its encoder block described in Appendix A.
The agent’s task is to reach a ball, which is food, that is located in the same room, which is 10 by 10 meters. Start agent and food positions are the same between episodes. All the agent needs to do is find the shortest path to the goal, as in the Gridworld experiments. In each episode, we measure the number of actions (steps) the agent takes to reach the goal.
The results are shown in Figure 5. We can see that after 150 episodes with the maximum number of steps set to 40, the agent stably reaches the goal. Although slight instability persists, the agent’s path is near optimal.
We weren’t able to train CSCG and LSTM with the same encoder block as we used for DHTM, even after increasing their trajectory buffer sizes. We think that the problem might be because of the non-stationarity of the encoder’s output, which adapts continuously.
Figure 5: Amount of action steps the agent uses to reach the goal in a 10x10-meter room in AnimalAI environment using DHTM memory for SF formation. The agent and the goal have the same initial positions throughout the episodes. Each timestep, the agent gets an image of a first-person view of the environment and produces one of three actions: turn left, turn right, or go forward.

### v3

@@ 3.2 Neural Implementation
Initially, cells have no segments. As learning progresses, new non-zero connections, grouped into segments, are created. In equation 4, we benefit from having sparse factor value vectors because their complexity depends linearly on the number of non-zero components. And that’s usually the case in our model due to the one-step Monte Carlo learning and the special form of the emission factors $F^{k}_{e}$:
$F^{k}_{e}(h^{k}_{t},\varphi^{k}_{t})=\mathbb{I}[h^{k}_{t}\in\mathrm{col}(\varphi^{k}_{t})],$ (5)
where $\mathbb{I}$—indicator function, $\mathrm{col}(\varphi^{k}_{t})$ is a set of hidden states connected to the feature state $\varphi^{k}_{t}$ forming a column. The shape of the emission factor is inspired by the presumed columnar structure of the neocortex and has been shown to induce a sparse transition matrix in HMM [39].
The segment likelihood $L_{l}$ resulting from the sum-product algorithm is computed as if the presynaptic cells were independent. However, this isn’t usually the case for sparse factors. To account for their interdependence, we substitute the following equation for segment log-likelihood:
$\log L_{l}$ $=w_{l}\log\frac{1}{n_{l}}\sum_{u\in\mathrm{rec}(l)}w_{ul}m_{u}+\sum_{u\in\mathrm{rec}(l)}(1-w_{ul})\log m_{u},$ (6)
where $w_{ul}$—synapse efficacy or neuron specificity for segment, such that ${w_{ul}=p(s_{l}=1|c^{u}_{t-1}=1)}$, $w_{l}=\frac{1}{n_{l}}\sum_{u}w_{ul}$—average synapse efficacy for segment $l$, and $n_{l}$-number of cells in segment’s receptive field.
The idea behind the formula is to approximate between two extreme cases. The first is ${p(s_{l}=1|c^{u}_{t-1}=1)}\to 1$ for all $u$, which means that all cells in the receptive field are dependent and part of a cluster, i.e. they fire together. In this case, ${p(s_{l})=m_{u}}$ for all $u$, but we also reduce the prediction variance by averaging across different $u$. In the opposite case, ${p(s_{l}=1|c^{u}_{t-1}=1)\to 0}$ for all $u$, presynaptic cells don’t cluster and the segment activation probability is just a product of the activation probability of each cell.
@@
where ${E_{l}=\log f_{l}+\log L_{l}}$, $\mathrm{cells}[H^{k}_{t}]$—indexes of cells representing states for the variable $H^{k}_{t}$. Here we also approximate the logarithmic sum with the $\mathrm{max}$ operation, inspired by the neurophysiological model of segment aggregation by cells [45].
The next step after computing the $p(h^{k}_{t})$ distribution parameters is to incorporate information about the current feature states ${p(h^{k}_{t}\ |\ \varphi^{k}_{t})\propto p(h^{k}_{t})\mathbb{I}[h^{k}_{t}\in\mathrm{col}(\varphi^{k}_{t})]}$. After that, the learning step is executed. The loop closing step of our TM algorithm is to assign the posterior of the current step ${p(h^{k}_{t}\ |\ \varphi^{k}_{t})}$ to $M^{i}_{t-1}$.
3.3 Learning
DHTM learns $f_{l}$ and $w_{ul}$ weights by Monte-Carlo Hebbian-like updates. First, $h^{i}_{t-1}$ and $h^{k}_{t}$ are sampled from their posterior distributions: ${p(h^{i}_{t-1}\ |\ \varphi^{i}_{t-1})\propto M^{i}_{t-1}}$ and ${p(h^{k}_{t}\ |\ \varphi^{k}_{t})}$ respectively. Then $f_{l}$ is updated according to the activity $s_{l}$ of the segment and the activity $c^{j}_{t}$ of its cell, so that $f_{l}$ is proportional to several coincidences ${s_{l}=c^{j}_{t}=1}$ in the recent past:
@@ 3.3 Learning
where $\alpha\in[0,1)$ is the segment’s learning rate.
That is, $f_{l}$ is trained to be equal to the average cell activation frequency when the segment is active. It’s similar to Baum-Welch’s update rule [41] for the transition matrix in HMM, which in effect counts transitions from one state to another, but in our case, the previous state (context) is represented by a group of RVs.
Weights $w_{ul}$ are also updated by the Hebbian rule to reflect the specificity of a presynaptic $u$ for activating a segment $l$. That is, they are targeted to represent the probability ${p(s_{l}=1\ |\ c^{u}_{t-1}=1)}$ that segment $s_{l}$ is active, given that cell $u$ was active at the previous time step. We could learn this by counting activation matches and mismatches. But in our algorithm, it is approximated as an exponential moving average of the frequency activation of segment $s_{l}$, given $c^{u}_{t-1}=1$:
$\Delta w_{ul}=\beta\cdot\mathbb{I}[c^{u}_{t-1}=1]\cdot(\mathbb{I}[s_{l}=1]-w_{ul}),$ (9)
where $\beta\in[0,1)$ — learning rate.
@@ 3.4 Agent Architecture
We incorporate DHTM as a part of an RL agent. The same agent is used for other memories tested. The agent consists of a memory model and a feature reward function. The memory model generates SF by predicting cumulative future distributions of feature variables $\Phi^{k}$. We assume that the reward function can be decomposed linearly in feature space as ${R_{t}=\frac{1}{n}\sum^{n}_{k=1}{r^{k}_{i}}}$, where $r^{k}_{i}$ is a reward associated with feature state $\varphi^{k}_{t}=i$ and $n$—number of feature variables. $r^{k}_{i}$ is learned during interaction with the environment and is used in combination with SF representations to estimate the action value function (see Appendix A for details).
@@
The agent training procedure is outlined in Algorithm 1. For each episode, the memory state is reset to a fixed initial message with RESET_MEMORY() and the variable action is set to a fixed initial action. An observation returned by an environment (obs) is encoded as a set of categorical variables. In the OBSERVE() routine, memory learns to predict next feature states as described in section 3. An agent learns associations to feature states and rewards in the REINFORCE() function:
@@
We include actions in DHTM by forcing some of the hidden variables $H^{k}_{t}$ to represent actions. That is, we assume that information about action is contained in the hidden state of the model. For example, if we have 4 actions, we set 4 states for one of the hidden variables and set its state from observing the action.
In SAMPLE_ACION(), for each action and current state, TM predicts feature states up to time step $T$ under policy $\pi$, and the predictions are combined to form SF (equation 11). For simplicity, we always build SF under a uniform policy, which doesn’t guarantee optimality, but we found that it doesn’t make a significant difference to iterative policy improvement in our experimental setups. The planning horizon $T$ is determined dynamically from the distribution of $\Phi^{k}_{t+l}$. If the distribution of $\Phi^{k}_{t+l}$ is close to uniform or states with positive reward have high probability, the prediction cycle is stopped. The action value is determined by combining the corresponding SF and feature state rewards as shown in equation 12. Finally, the action is sampled from the softmax distribution over the action values.
$\mathrm{SF}^{\pi}_{t+T}(\varphi^{k}=j\ |\ h_{t})=\sum^{T}_{l=0}\gamma^{l}p_{\pi}(\varphi^{k}_{t+l+1}=j\ |\ h_{t})$ (11)
$Q^{\pi}(h_{t},a_{t})=\sum_{jk}\sum_{h_{t+1}}\mathrm{SF}^{\pi}_{t+T}(\varphi^{k}=j\ |\ h_{t+1})\cdot p(h_{t+1}\ |\ h_{t},a_{t})\cdot r^{k}_{j},$ (12)
3.5 DHTM and Episodic Control
Algorithm 2 Episodic memory learning
0:  $\varphi_{t+1}$, $a_{t}$
1:  $h_{t+1}$leftarrow $\mathcal{D}_{a_{t}}(h_{t})$
[…]
3:  if $h_{t+1}$ is null or $\varphi^{*}_{t+1}$ is not $\varphi_{t+1}$ then
4:     $h_{t+1}$leftarrow $\underset{h\not\in K}{\mathrm{argmax}\ }F_{e}(h,\varphi_{t+1})$ $K$ is the set of keys for all actions
5:     $\mathcal{D}_{a_{t}}(h_{t})$leftarrow $h_{t+1}$
6:  end if
7:  $h_{t}$leftarrow $h_{t+1}$
It can be shown that DHTM combined with the agent architecture described in the previous section is similar to episodic control.
DTHM performs best when its hidden state space is large. Since the hidden state for unpredictable observations is formed randomly, it’s important that there are no accidental state collisions when remembering transitions. That is, DHTM tries to remember new sequences without interfering with old memories. In effect, this makes DHTM a buffer of agent trajectories with a built-in recall engine. Episodic Control is based on a similar idea, and to investigate its difference from DHTM, we implemented an Episodic Control agent, a simplified counterpart of DHTM, reflecting its core property that presumably helps it solve RL tasks.
In our EC version, $K_{a}$ is an array of hidden states $h_{t}$ and $V_{a}$ is an array of corresponding subsequent hidden states $h_{t+1}$. If the key state $h_{t}$ already exists in the dictionary, its value is rewritten. That is, $\mathcal{D}_{a}$ represents the transition matrix of a deterministic HMM for a feature variable $\Phi_{t}$. The hidden state $h_{t}$ and the feature state $\varphi_{t}$ have the same dependence as in DHTM (equation 5), but here we assume infinite cells per column. To predict the next state, we just have to look up $\mathcal{D}_{a}$ for the current state $h_{t}$, but if there is no entry for $h_{t}$ or the prediction does not match the observed state $\varphi_{t+1}$, then the new state is formed. Unlike DHTM, the new state is not completely random, but to avoid collisions, the algorithm makes sure that this state hasn’t been chosen before. The learning procedure of the EC agent’s memory $\mathcal{D}_{a}$ is summarized in the algorithm 2.
Since $\mathcal{D}_{a}$ can be transformed into a transition matrix, the process of forming SF can be reduced to one in Sec. 3.4. However, since the resulting transition matrix is sparse, it’s more efficient to consider $\mathcal{D}_{a}$ as a graph representation and to perform SF formation as a breadth-first search (see Appendix D.3 for details).
@@ 4 Experiments
We test our model in a reinforcement learning task in the Gridworld environment and in a more challenging AnimalAI 3D environment. This section shows how different memory models affect the adaptability of an RL agent. For each setup, the results are averaged over five experiments with different seed values, and the standard deviation is shown. See Appendix D for details on the baselines, their model parameters, and training regimes. See Appendix E for a more detailed description of the environments and setups. The source code of the experiments and algorithms is at https://anonymous.4open.science/r/dhtm-FA0E.
@@ 4.1 Gridworld
Figure 2: POMDP 5x5 Gridworld foraging task using different types of memory: DHTM, CSCG, LSTM. CSCG and LSTM are trained in two modes. The first mode is with oracle, which tells the agent to reinitialize memory and clear the trajectory buffer during environmental change. The second mode is without oracle, so the agent continues using its current strategy to gather new trajectories and update memory. There is also an episodic control agent (ec) among baselines, which uses a simple dictionary to store trajectories and make predictions. A. Dynamic of steps required for the agent to reach the goal. In the 300th episode, the goal is blocked by a wall, so the agent needs to find a new optimal trajectory. B. Total steps taken during 1000 episodes.
The first test was conducted in partially observable Gridworld in a 5x5 alternating maze with a stationary end state and the agent’s starting position. The agent can observe the colour of the cell it is in and the colour of an obstacle it encounters, while being punished in addition to the constant step penalty. Since the observation can be represented directly as a categorical variable with the number of states equal to the number of possible colours, no encoder is needed. Each episode starts with the agent in the same position and ends when the agent enters the terminal state (i.e., when the reward is collected) or when the maximum number of action steps is reached. In each episode, we measure the number of action steps required for the agent to reach the goal.
Each trial lasts $1000$ episodes and consists of two phases. In the first phase, which lasts $300$ episodes, there are no obstacles on the shortest path to the terminal state. In the second phase, the previously optimal path is blocked by a wall, so the agent has to take a detour to get the reward. The experiment aims to show the ability of a memory to relearn the structure of the maze in the test time. If the agent is able to reach the end state in the second phase, it suggests that the model can forget, or at least suppress, the memory of the old shortest path when it is blocked. Otherwise, the agent will continue to hit the wall and fail to reach the terminal state.
In this experiment, we compare our memory model DHTM with CSCG [39] and LSTM [9] within the same agent architecture that uses memory to form successor features and evaluate actions (see Sec. 3.4). CSCG and LSTM are trained in two modes. The first mode is with an oracle that tells the agent to reinitialize memory and clear the trajectory buffer just in time for the environmental change. The second mode is without an oracle, so the agent continues to use its current strategy to collect new trajectories and update memory in the second phase. There is also a simple episodic control agent (EC) among the baselines, which uses a dictionary to store trajectories and make predictions (see Section 3.5 for details).
The results in Figure 2 show that CSCG effectively can’t adapt to the environmental change in the online regime, but only if it is retrained from scratch on new trajectories. And, apparently, LSTM is only able to learn a predictive model for a small horizon, since it performs significantly worse than other baselines in the second phase, even with oracle enabled. In contrast, DHTM and EC are able to adapt online without explicit memory reset. Another important observation is that the EC agent performs on par with DHTM, suggesting that its mechanism can explain most of DHTM’s performance in this task.
@@ 4.2 AnimalAI
Figure 3: Amount of action steps the agent performs to reach the goal in a 10x10-meter room in AnimalAI environment using DHTM, CSCG, LSTM and EC dictionary. The agent and the goal have the same initial positions throughout the episodes. Each time-step, the agent gets an encoded image of a first-person view of the environment and produces one of three actions: turn left, turn right, or go forward. A. Dynamic of steps required for the agent to reach the goal. B. Total steps taken during 1000 episodes by the agent with DHTM with K-Means encoder and Categorical VAE. For the latter, DHTM’s context factor size is varied (i.e., how many previous time-step hidden variables are considered for prediction).
AnimalAI is a 3D environment capable of providing rich visual input to an agent in the form of RGB images. The goal of the experiment described here is to show that DHTM can handle distributed representations and is scalable to environments with larger state spaces.
[…]
We test the agent using two different encoders: K-Means and Categorical VAE [46]. Both encoders are pre-trained on RGB images collected with a uniform strategy during $2000$ episodes in the same room used for agent testing. K-Means represents an image as a single categorical variable with $50$ states, and Categorical VAE is configured for distributed representation with five variables with $50$ states each (see Appendix C for details).
[…]
To show that DHTM can’t simply be replaced by multiple parallel EC dictionaries or CSCGs in the case of distributed encoding, we also vary the context factor size of DHTM, that is, how many previous time-step variables are taken into account to predict the current state of a variable. The factor size of 1 corresponds to the graph of the Factorial HMM, with parallel HMMs. It can be seen from Figure 3B that in our case of categorical VAE, apparently knowing only the previous state of the variable is not enough to reliably predict rewarding features, since the agent doesn’t converge to the shortest path. We also tested the variant (factor 1*) where each feature variable is represented by three hidden states, to ensure that the result for factor size $1$ is not due to DHTM’s hidden state collisions. DHTM with a factor that takes into account more than one previous time-step hidden variable (excluding the action variable) performs significantly better. For diagrams depicting factor graphs used in these experiments, see Appendix D.4.